OpenStudy (anonymous):

Water coming out from a fountain is modeled by the function f(x) = -x2 + 6x + 6 where f(x) represents the height, in feet, of the water from the fountain at different times x, in seconds. What does the average rate of change of f(x) from x = 2 to x = 5 represent?

OpenStudy (anonymous):

The answer choices are: The water travels an average distance of 4 feet from 2 seconds to 5 seconds. The water travels an average distance of 1 foot from 2 seconds to 5 seconds. The water falls down with an average speed of 2 feet per second from 2 seconds to 5 seconds. The water falls down with an average speed of 1 foot per second from 2 seconds to 5 seconds.

OpenStudy (anonymous):

Do you know how to do this?

OpenStudy (anonymous):

I don't want to just give you the answer.

OpenStudy (anonymous):

average rate of change = (f(5) - f(2))/(5 - 2) = (-25 + 30 + 6 - (-4 + 12 + 6))/3 = (11 - 14)/3 = -1

OpenStudy (anonymous):

You googled that you butt.

OpenStudy (anonymous):

dang it hotchellerae i was trying to answer it myself

OpenStudy (anonymous):

But I believe it's D, since you can already see the work.

OpenStudy (anonymous):

sorry ok i will delete it

OpenStudy (anonymous):

thanks guys ily!!!!!!